OpenStudy (anonymous):

working out to 4(2-b) + 3(3+2b) = 33

OpenStudy (anonymous):

8 - 4b + 9 + 6b = 33 therefore 2b = 16 and b = 8

OpenStudy (anonymous):

where did you get 6 from on 2b = 16 ?

OpenStudy (anonymous):

16*

OpenStudy (anonymous):

Multiply 4(2-b) giving you 8-4b, then multiply 3(3+2b) giving you 9+6b. The equation will then be 8-4b+9+6b=33 after this step, take the 9 and 9 and subtract from the left fside of the equation to the right, giving you -4b+6b=33-9-8. That will give you -2b = 16

OpenStudy (anonymous):

Sorry 2b=16

OpenStudy (anonymous):

b=8